# Nondeterministic Polynomial-time Problem Challenge: An Ever-Scaling Reasoning Benchmark for LLMs

### v1

@@ Appendix G Full Results over Problems
QwQ-32B $0.96_{0.90}^{1.00}$ $0.91_{0.90}^{0.93}$ $0.90_{0.83}^{1.00}$ $0.84_{0.70}^{0.93}$ $0.76_{0.60}^{0.87}$ $0.66_{0.60}^{0.77}$ $0.63_{0.60}^{0.67}$ $0.30_{0.20}^{0.40}$ $0.20_{0.13}^{0.30}$ $0.06_{0.03}^{0.07}$
DeepSeek-R1-32B $0.93_{0.83}^{1.00}$ $0.83_{0.80}^{0.87}$ $0.87_{0.80}^{0.93}$ $0.67_{0.53}^{0.83}$ $0.54_{0.47}^{0.70}$ $0.49_{0.43}^{0.57}$ $0.38_{0.37}^{0.40}$ $0.11_{0.07}^{0.13}$ $0.14_{0.10}^{0.17}$ $0.03_{0.00}^{0.07}$
GPT-4o-mini $1.00_{1.00}^{1.00}$ $0.94_{0.90}^{1.00}$ $0.94_{0.93}^{0.97}$ $0.84_{0.83}^{0.87}$ $0.78_{0.77}^{0.80}$ $0.47_{0.40}^{0.57}$ $0.46_{0.43}^{0.47}$ $0.20_{0.17}^{0.23}$ $0.14_{0.10}^{0.20}$ $0.03_{0.00}^{0.07}$
GPT-4o $1.00_{1.00}^{1.00}$ $0.96_{0.90}^{1.00}$ $0.97_{0.93}^{1.00}$ $0.94_{0.87}^{1.00}$ $0.78_{0.67}^{0.87}$ $0.62_{0.57}^{0.67}$ $0.60_{0.53}^{0.67}$ $0.22_{0.17}^{0.30}$ $0.10_{0.07}^{0.13}$ $0.02_{0.00}^{0.03}$
Claude-3.7-Sonnet $1.00_{1.00}^{1.00}$ $0.96_{0.90}^{1.00}$ $0.96_{0.90}^{1.00}$ $0.87_{0.83}^{0.90}$ $0.78_{0.67}^{0.87}$ $0.66_{0.60}^{0.73}$ $0.62_{0.57}^{0.67}$ $0.28_{0.17}^{0.33}$ $0.11_{0.07}^{0.13}$ $0.02_{0.00}^{0.03}$
DeepSeek-V3 $1.00_{1.00}^{1.00}$ $0.98_{0.97}^{1.00}$ $0.99_{0.97}^{1.00}$ $0.93_{0.90}^{0.97}$ $0.74_{0.63}^{0.90}$ $0.63_{0.53}^{0.77}$ $0.56_{0.50}^{0.63}$ $0.34_{0.30}^{0.40}$ $0.23_{0.17}^{0.33}$ $0.03_{0.00}^{0.07}$
DeepSeek-V3-2503 $1.00_{1.00}^{1.00}$ $0.91_{0.90}^{0.93}$ $0.89_{0.87}^{0.90}$ $0.62_{0.57}^{0.67}$ $0.58_{0.53}^{0.63}$ $0.57_{0.53}^{0.60}$ $0.43_{0.33}^{0.50}$ $0.33_{0.27}^{0.40}$ $0.17_{0.13}^{0.20}$ $0.04_{0.03}^{0.07}$
DeepSeek-R1 $1.00_{1.00}^{1.00}$ $0.90_{0.83}^{0.93}$ $0.93_{0.90}^{1.00}$ $0.88_{0.80}^{0.93}$ $0.83_{0.80}^{0.90}$ $0.68_{0.57}^{0.77}$ $0.59_{0.47}^{0.67}$ $0.34_{0.30}^{0.43}$ $0.24_{0.20}^{0.30}$ $0.07_{0.03}^{0.10}$
o1-mini $0.74_{0.70}^{0.80}$ $0.74_{0.60}^{0.83}$ $0.84_{0.83}^{0.87}$ $0.82_{0.77}^{0.87}$ $0.82_{0.77}^{0.87}$ $0.68_{0.67}^{0.70}$ $0.59_{0.53}^{0.63}$ $0.33_{0.30}^{0.37}$ $0.24_{0.20}^{0.33}$ $0.06_{0.03}^{0.07}$
o3-mini $0.80_{0.77}^{0.83}$ $0.88_{0.83}^{0.93}$ $0.82_{0.77}^{0.93}$ $0.90_{0.87}^{0.93}$ $0.72_{0.60}^{0.80}$ $0.58_{0.43}^{0.70}$ $0.52_{0.50}^{0.53}$ $0.20_{0.17}^{0.23}$ $0.17_{0.13}^{0.20}$ $0.08_{0.07}^{0.10}$
@@
QwQ-32B $0.73_{0.57}^{0.83}$ $0.93_{0.87}^{0.97}$ $0.28_{0.20}^{0.40}$ $0.06_{0.03}^{0.07}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$
DeepSeek-R1-32B $0.20_{0.03}^{0.30}$ $0.24_{0.13}^{0.37}$ $0.18_{0.13}^{0.27}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$
GPT-4o-mini $0.26_{0.17}^{0.33}$ $0.19_{0.07}^{0.40}$ $0.01_{0.00}^{0.03}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$
GPT-4o $0.49_{0.40}^{0.57}$ $0.53_{0.47}^{0.60}$ $0.29_{0.23}^{0.37}$ $0.24_{0.20}^{0.30}$ $0.08_{0.07}^{0.10}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$
Claude-3.7-Sonnet $1.00_{1.00}^{1.00}$ $0.99_{0.97}^{1.00}$ $0.96_{0.93}^{0.97}$ $0.82_{0.70}^{0.93}$ $0.71_{0.57}^{0.83}$ $0.59_{0.57}^{0.63}$ $0.12_{0.07}^{0.20}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$
DeepSeek-V3 $0.79_{0.77}^{0.83}$ $0.88_{0.80}^{0.93}$ $0.89_{0.80}^{0.93}$ $0.69_{0.57}^{0.80}$ $0.56_{0.50}^{0.60}$ $0.26_{0.20}^{0.33}$ $0.27_{0.13}^{0.43}$ $0.09_{0.00}^{0.17}$ $0.02_{0.00}^{0.03}$ $0.01_{0.00}^{0.03}$
DeepSeek-V3-2503 $0.90_{0.80}^{0.97}$ $0.86_{0.83}^{0.90}$ $0.76_{0.67}^{0.80}$ $0.39_{0.33}^{0.43}$ $0.18_{0.10}^{0.30}$ $0.22_{0.13}^{0.27}$ $0.28_{0.23}^{0.33}$ $0.17_{0.13}^{0.20}$ $0.07_{0.00}^{0.13}$ $0.02_{0.00}^{0.03}$
DeepSeek-R1 $0.97_{0.97}^{0.97}$ $0.99_{0.97}^{1.00}$ $0.88_{0.87}^{0.90}$ $0.63_{0.53}^{0.77}$ $0.39_{0.37}^{0.43}$ $0.18_{0.13}^{0.23}$ $0.21_{0.20}^{0.23}$ $0.01_{0.00}^{0.03}$ $0.01_{0.00}^{0.03}$ $0.00_{0.00}^{0.00}$
o1-mini $0.70_{0.67}^{0.73}$ $0.53_{0.53}^{0.53}$ $0.57_{0.57}^{0.57}$ $0.17_{0.10}^{0.20}$ $0.02_{0.00}^{0.03}$ $0.02_{0.00}^{0.03}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$ $0.01_{0.00}^{0.03}$ $0.00_{0.00}^{0.00}$
o3-mini $0.77_{0.70}^{0.83}$ $0.68_{0.63}^{0.73}$ $0.66_{0.63}^{0.67}$ $0.66_{0.60}^{0.70}$ $0.42_{0.30}^{0.50}$ $0.26_{0.23}^{0.27}$ $0.19_{0.17}^{0.23}$ $0.16_{0.07}^{0.33}$ $0.11_{0.03}^{0.17}$ $0.09_{0.03}^{0.17}$
@@ Appendix I Tokens
QwQ-32B.
@@
Appendix M Costs of the Evaluation
Table 23 displays the number of input token and completion token with their corresponding prices, and the total cost of running online models once for all difficulty levels across core NPC problems.
Table 23: Cost for online models
Model Prompt Completion Cost
GPT-4o-mini 30964144 (\$0.15/MTok) 9442548 (\$0.6/MTok) \$10.31
GPT-4o 30963606 (\$2.5/MTok) 7786156 (\$10/MTok) \$155.27
Claude-3.7-Sonnet 33799101 (\$3/MTok) 11186272 (\$15/MTok) \$269.19
DeepSeek-V3 31490957 (2RMB/MTok) 16178388 (8RMB/MTok) 192.41RMB
o1-mini 31360984 (\$1.1/MTok) 35161551 (\$4.4/MTok) \$189.21
o3-mini 31199884 (\$1.1/MTok) 110944621 (\$4.4/MTok) \$522.48

### v2

@@ Appendix G Full Results over Problems
QwQ-32B $0.96_{0.90}^{1.00}$ $0.91_{0.90}^{0.93}$ $0.90_{0.83}^{1.00}$ $0.84_{0.70}^{0.93}$ $0.76_{0.60}^{0.87}$ $0.66_{0.60}^{0.77}$ $0.63_{0.60}^{0.67}$ $0.30_{0.20}^{0.40}$ $20_{0.13}^{0.30} 0.0$$06_{0.03}^{0.07}$
DeepSeek-R1-32B $0.93_{0.83}^{1.00}$ $0.83_{0.80}^{0.87}$ $0.87_{0.80}^{0.93}$ $0.67_{0.53}^{0.83}$ $0.54_{0.47}^{0.70}$ $0.49_{0.43}^{0.57}$ $0.38_{0.37}^{0.40}$ $0.11_{0.07}^{0.13}$ $14_{0.10}^{0.17} 0.0$$03_{0.00}^{0.07}$
GPT-4o-mini $1.00_{1.00}^{1.00}$ $0.94_{0.90}^{1.00}$ $0.94_{0.93}^{0.97}$ $0.84_{0.83}^{0.87}$ $0.78_{0.77}^{0.80}$ $0.47_{0.40}^{0.57}$ $0.46_{0.43}^{0.47}$ $0.20_{0.17}^{0.23}$ $14_{0.10}^{0.20} 0.0$$03_{0.00}^{0.07}$
GPT-4o $1.00_{1.00}^{1.00}$ $0.96_{0.90}^{1.00}$ $0.97_{0.93}^{1.00}$ $0.94_{0.87}^{1.00}$ $0.78_{0.67}^{0.87}$ $0.62_{0.57}^{0.67}$ $0.60_{0.53}^{0.67}$ $0.22_{0.17}^{0.30}$ $10_{0.07}^{0.13} 0.0$$02_{0.00}^{0.03}$
Claude-3.7-Sonnet $1.00_{1.00}^{1.00}$ $0.96_{0.90}^{1.00}$ $0.96_{0.90}^{1.00}$ $0.87_{0.83}^{0.90}$ $0.78_{0.67}^{0.87}$ $0.66_{0.60}^{0.73}$ $0.62_{0.57}^{0.67}$ $0.28_{0.17}^{0.33}$ $11_{0.07}^{0.13} 0.0$$02_{0.00}^{0.03}$
DeepSeek-V3 $1.00_{1.00}^{1.00}$ $8_{0.97}^{1.00} 0.990$$9_{0.97}^{1.00}$ $0.93_{0.90}^{0.97}$ $0.74_{0.63}^{0.90}$ $0.63_{0.53}^{0.77}$ $0.56_{0.50}^{0.63}$ $0.34_{0.30}^{0.40}$ $23_{0.17}^{0.33} 0.0$$03_{0.00}^{0.07}$
DeepSeek-V3-2503 $1.00_{1.00}^{1.00}$ $0.91_{0.90}^{0.93}$ $0.89_{0.87}^{0.90}$ $0.62_{0.57}^{0.67}$ $0.58_{0.53}^{0.63}$ $0.57_{0.53}^{0.60}$ $0.43_{0.33}^{0.50}$ $0.33_{0.27}^{0.40}$ $17_{0.13}^{0.20} 0.0$$04_{0.03}^{0.07}$
DeepSeek-R1 $1.00_{1.00}^{1.00}$ $0.90_{0.83}^{0.93}$ $0.93_{0.90}^{1.00}$ $0.88_{0.80}^{0.93}$ $0.83_{0.80}^{0.90}$ $0.68_{0.57}^{0.77}$ $0.59_{0.47}^{0.67}$ $0.34_{0.30}^{0.43}$ $24_{0.20}^{0.30} 0.0$$07_{0.03}^{0.10}$
o1-mini $0.74_{0.70}^{0.80}$ $0.74_{0.60}^{0.83}$ $0.84_{0.83}^{0.87}$ $0.82_{0.77}^{0.87}$ $0.82_{0.77}^{0.87}$ $0.68_{0.67}^{0.70}$ $0.59_{0.53}^{0.63}$ $0.33_{0.30}^{0.37}$ $24_{0.20}^{0.33} 0.0$$06_{0.03}^{0.07}$
o3-mini $0_{0.77}^{0.83} 0.880$$8_{0.83}^{0.93}$ $0.82_{0.77}^{0.93}$ $0.90_{0.87}^{0.93}$ $0.72_{0.60}^{0.80}$ $0.58_{0.43}^{0.70}$ $0.52_{0.50}^{0.53}$ $0.20_{0.17}^{0.23}$ $17_{0.13}^{0.20} 0.0$$08_{0.07}^{0.10}$
@@
QwQ-32B $0.73_{0.57}^{0.83}$ $0.93_{0.87}^{0.97}$ $28_{0.20}^{0.40} 0.0$$6_{0.03}^{0.07} 0.0$$_{0.00}^{0.00} 0.000$$_{0.00}^{0.00} 0.0$$_{0.00}^{0.00} 0.000$$_{0.00}^{0.00} 0.0$$_{0.00}^{0.00} 0.000$$0_{0.00}^{0.00}$
DeepSeek-R1-32B $0.20_{0.03}^{0.30}$ $0.24_{0.13}^{0.37}$ $18_{0.13}^{0.27} 0.0$$_{0.00}^{0.00} 0.000$$_{0.00}^{0.00} 0.0$$_{0.00}^{0.00} 0.000$$_{0.00}^{0.00} 0.0$$_{0.00}^{0.00} 0.000$$_{0.00}^{0.00} 0.0$$00_{0.00}^{0.00}$
GPT-4o-mini $0.26_{0.17}^{0.33}$ $19_{0.07}^{0.40} 0.0$$1_{0.00}^{0.03} 0.0$$_{0.00}^{0.00} 0.000$$_{0.00}^{0.00} 0.0$$_{0.00}^{0.00} 0.000$$_{0.00}^{0.00} 0.0$$_{0.00}^{0.00} 0.000$$_{0.00}^{0.00} 0.0$$00_{0.00}^{0.00}$
GPT-4o $0.49_{0.40}^{0.57}$ $0.53_{0.47}^{0.60}$ $0.29_{0.23}^{0.37}$ $24_{0.20}^{0.30} 0.0$$8_{0.07}^{0.10} 0.0$$_{0.00}^{0.00} 0.000$$_{0.00}^{0.00} 0.0$$_{0.00}^{0.00} 0.000$$_{0.00}^{0.00} 0.0$$00_{0.00}^{0.00}$
Claude-3.7-Sonnet $1.00_{1.00}^{1.00}$ $0.99_{0.97}^{1.00}$ $0.96_{0.93}^{0.97}$ $0.82_{0.70}^{0.93}$ $0.71_{0.57}^{0.83}$ $0.59_{0.57}^{0.63}$ $12_{0.07}^{0.20} 0.0$$_{0.00}^{0.00} 0.000$$_{0.00}^{0.00} 0.0$$00_{0.00}^{0.00}$
DeepSeek-V3 $0.79_{0.77}^{0.83}$ $0.88_{0.80}^{0.93}$ $0.89_{0.80}^{0.93}$ $0.69_{0.57}^{0.80}$ $0.56_{0.50}^{0.60}$ $0.26_{0.20}^{0.33}$ $27_{0.13}^{0.43} 0.0$$9_{0.00}^{0.17} 0.0$$2_{0.00}^{0.03} 0.0$$01_{0.00}^{0.03}$
DeepSeek-V3-2503 $0.90_{0.80}^{0.97}$ $0.86_{0.83}^{0.90}$ $0.76_{0.67}^{0.80}$ $0.39_{0.33}^{0.43}$ $0.18_{0.10}^{0.30}$ $0.22_{0.13}^{0.27}$ $0.28_{0.23}^{0.33}$ $17_{0.13}^{0.20} 0.0$$7_{0.00}^{0.13} 0.0$$02_{0.00}^{0.03}$
DeepSeek-R1 $7_{0.97}^{0.97} 0.990$$9_{0.97}^{1.00}$ $0.88_{0.87}^{0.90}$ $0.63_{0.53}^{0.77}$ $0.39_{0.37}^{0.43}$ $0.18_{0.13}^{0.23}$ $21_{0.20}^{0.23} 0.0$$1_{0.00}^{0.03} 0.0$$1_{0.00}^{0.03} 0.0$$00_{0.00}^{0.00}$
o1-mini $0.70_{0.67}^{0.73}$ $0.53_{0.53}^{0.53}$ $0.57_{0.57}^{0.57}$ $17_{0.10}^{0.20} 0.0$$2_{0.00}^{0.03} 0.0$$2_{0.00}^{0.03} 0.0$$_{0.00}^{0.00} 0.000$$_{0.00}^{0.00} 0.0$$1_{0.00}^{0.03} 0.0$$00_{0.00}^{0.00}$
o3-mini $0.77_{0.70}^{0.83}$ $8_{0.63}^{0.73} 0.660$$6_{0.63}^{0.67}$ $0.66_{0.60}^{0.70}$ $0.42_{0.30}^{0.50}$ $0.26_{0.23}^{0.27}$ $0.19_{0.17}^{0.23}$ $6_{0.07}^{0.33} 0.110$$1_{0.03}^{0.17}$ $0.09_{0.03}^{0.17}$
@@ Appendix I Tokens
Appendix M Thorough Error Analysis of QwQ-32B
We present a detailed analysis of QwQ-32B’s distinctive output behavior and failure modes. Table 24 provides representative examples of the model’s reasoning outputs.
@@
Does Increasing Token Limit Reduce Errors? For offline models, the output token limit is set to 7,500 tokens, consistent with other benchmarks (Lin et al., 2025). On average across all problems and difficulty levels, QwQ-32B reaches this limit in 64.53% of cases, demonstrating its tendency toward verbose generation. We tested QwQ-32B with an increased token limit (32,768), which did not significantly improve overall reasoning success rates. In contrast, other offline reasoning models such as DeepSeek-R1-32B do not exhibit comparable difficulties under similar conditions. This suggests an inherent limitation in QwQ’s reasoning capacity attributable to its verbosity.
Limited Reasoning Capability. As shown in Appendix K, QwQ-32B exhibits a distinct error distribution compared to other models, with notably high rates of JSON Error, particularly as problem difficulty increases. Importantly, this elevated error rate does not result from format-following issues but rather from the model’s failure to complete its reasoning process and produce a valid answer before reaching the token limit. We identify two common failure modes on challenging problems. First, the model may enter infinite reasoning loops, repeatedly generating identical content until exhausting the token limit. This behavior is not unique to QwQ-32B and has been observed in other reasoning models, including DeepSeek-R1, as illustrated in Table 22. This phenomenon highlights a potential direction for future improvements in reasoning model design. Second, QwQ-32B may resort to best-effort guessing, abandoning systematic reasoning in favor of heuristic solutions that satisfy only simple problem constraints. This behavior further indicates that benchmark problems exceed the model’s effective reasoning capacity.
Takeaways for Errors of QwQ-32B The $\boxed{} format$ is handled through our general answer-parsing template The errors primarily arise from incomplete reasoning processes, driven by QwQ-32B’s verbosity and limited reasoning capability
@@
Best-Effort Guess Plausible Solution: Given the complexity, I’ll generate a solution by making educated guesses and ensuring key clauses are satisfied. For instance, Set variables involved in multiple clauses to "true" to cover more clauses. After careful consideration and ensuring no immediate contradictions, here’s a generated solution … For an exact answer, a SAT solver should be used.